# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ 3. Quantum Algorithms Complexity
Consider a random oracle. Yamakawa and Zhandry (yamakawa2022verifiable, ) proved unconditionally unless otherwise stated, that relative to the oracle, with probability one, it holds that:
@@
  • The results in (yamakawa2022verifiable, ), as stated by the authors, do not contradict the conjecture of Aaronson and Ambanis (aaronson2014need, ). As a conjecture, the authors mentioned that considering the Aaronson-Ambanis conjecture is true, one may conclude that there is publicly verifiable certifiable randomness with minimal rounds of interaction.
@@
The structure of optimally entangled states is constrained by quantum theory’s constant character (alber1999entanglement, ). Abrams et al. (abrams1998nonlinear, ) illustrated that one may exploit nonlinear time evolution to allow a quantum computer to solve NP-complete and #P problems in polynomial time. The authors showed how to accomplish this exponential speed-up using the Weinberg model of nonlinear quantum mechanics. All known experiments have confirmed the linearity of quantum mechanics to a high degree of accuracy. So, the result in (abrams1998nonlinear, ) may be seen as evidence.
@@
Wang (wang2022classically, ) leveraged existing classical techniques to improve quantum optimization. In particular, he first pre-processed the input with a classical algorithm, obtaining an approximate solution. Second, a quantum circuit to search the neighborhood of the solution of the classical algorithm, aiming for higher-quality solutions. The author introduced the Classically-Boosted Quantum Optimization Algorithm (CBQOA). One may solve a wide range of combinatorial optimization problems with CBQOA. E.g. all unconstrained problems, constrained problems such as Max Bisection, Maximum Independent Set, Minimum Vertex Cover, Portfolio Optimization, Traveling Salesperson, etc. A crucial component of CBQOA is a practice-friendly continuous-time quantum walk (CTQW) on a properly constructed graph, connecting the feasible solutions. CBQOA utilizes CTQW, the output of a classical procedure, constructing a superposition of feasible solutions. The algorithm solves constrained problems without applying changes to their cost functions. It then narrows down evaluating the quantum state to the feasible subspace. The algorithm also relies not on indexing the feasible solutions. Wang illustrates the applications of CBQOA to Max 3SAT, and Max Bisection as well. The author further provides empirical results that CBQOA outperforms previous approaches.
Jeronimo and Wu (10.1145/3564246.3585248, ) denoted the explored the class $\text{QMA}^{+}(2)$ as the unentangled quantum proofs with non-negative amplitudes. Under the above setting, one may design proof verification protocols for (increasingly) hard problems, the authors stated. This may be done via logarithmic size quantum proofs, and also by having a constant probability gap that allows distinguishing yes from no instances. Specifically, the authors designed global protocols for small set expansion, unique games, and PCP verification. Consequently, Jeronimo and Wu had the result of $\text{NP}\subseteq\text{QMA}^{+}_{\log}(2)$ with a constant gap. Taking advantage of the mentioned constant gap, the authors could scale up the aforementioned result to $\text{QMA}^{+}(2)$. So, they were able to obtain $\text{QMA}^{+}(2)=\text{NEXP}$ by setting up stronger in-detail properties of the PCP for NEXP.
The classical matching theory is a mathematical method for finding pairs between elements from two sets. One may optimize it according to certain criteria. Classical matching theory can be defined as matrices that do not contain negative entries. In quantum theory, the positive operator concept is a generalization of matrices that have nonnegative entries. In terms of complete positive operators, or equivalently with bipartite density matrices, Gurvits (GURVITS2004448, ) reformulate the Edmonds problem. Edmonds’ problem is the central problem in the field of algebraic complexity, which J. Edmonds presented in 1967. This problem aims to decide if the span of a given tuple of square matrices contains a non-singular matrix. In other words, it investigates whether a given linear subspace of a matrix has a full rank. One of the most important cases when Edmonds’ problem can be solved in polynomial deterministic time, i.e., an intersection of two geometric matroids, corresponds to unentangled (a.k.a. separable) bipartite density matrices.
Denote $M(N)$ as the linear space of complex $N\crossproduct N$ matrices. To solve Edmonds’ problem, the author in (GURVITS2004448, ) introduced a general class of linear subspaces $M(N)$. In this class, there is a polynomial deterministic time algorithm to solve Edmonds’ problem (GURVITS2004448, ). Finally, Gurvits proved that the weak membership problem for the convex set of separable normalized bipartite density matrices is NP-hard.
@@
Quantum Merlin Arthur (QMA) represents the set of languages for which, when a string is in the language, there is a polynomial-size quantum proof that convinces a polynomial time quantum verifier of this fact with high probability. In addition, the verifier with high probability rejects every linear size quantum state where no string is included in the language.
[…]
@@
One may generalize the satisfiability problem in classical computation to a constraint satisfaction problem in the quantum model, called Quantum Satisfiability. Consider a quantum k-SAT problem. Each constraint is then determined by a k-local projector. Also, a constraint is satisfied by all states in its null space. A previous work by Bravyi (bravyi2011efficient, ) showed that one may solve quantum 2-SAT on a classical computer. The author further proved that quantum k-SAT with $k\geq 4$ is QMA1-complete. Previously, Quantum 3-SAT was recognized as being in QMA1 (bravyi2011efficient, ). Nonetheless, its computational hardness was unknown. Gosset and Nagaj (gosset2016quantum, ) then proved that quantum 3-SAT is QMA1-hard. So, this problem is complete for this complexity class.
Khadiev and Safina (khadiev2023quantum, ) suggested a quantum algorithm for the dynamic programming approach for problems on directed acyclic graphs (DAGs) like $G(V,E)$. Denote $|V|=n$ as the number of vertices of $G(V,E)$. Further, let $\hat{n}$ be the number of vertices with at least one outgoing edge, and $|E|=m$ be the number of edges. Their algorithm’s running time is proportional to $\mathcal{O}(\sqrt{\hat{n}m}\log\hat{n})$. Moreover, executing the best-known deterministic algorithm up to the date of the publication of (khadiev2023quantum, ) takes $\mathcal{O}(n+m)$.
@@
The authors proved that their algorithm may solve problems that involve functions OR, AND, NAND, MAX, and MIN. For example, a Boolean formula represented by Zhegalkin polynomial (zhegalkin1927technique, ; gindikin1985algebraic, ; yablonsky1989introduction, ). Also, a Boolean circuit with shared input and non-constant depth evaluation. In addition, the single source longest paths search for weighted DAGs problem and the diameter search problem for unweighted DAGs.
@@
Quantum annealing algorithms are introduced as a solution to combinatorial optimization problems, by utilizing quantum fluctuations to escape local minima in energy landscapes, which able them to solve complex NP-hard problems. Cruz-Santos et al. (cruz2019qubo, ) aimed to apply quantum annealing to the theory of cuts, which is a crucial field in theoretical computer science. The authors developed a method to convert the Minimum Multicut Problem into the Quadratic Unconstrained Binary Optimization (QUBO) representation and have addressed the technical challenges that arise when submitting a problem to the quantum annealer processor. Two constructions of the quadratic unconstrained binary optimization functions for the Minimum Multicut Problem have been provided and compared the tradeoffs between the two mappings. Also, numerical scaling analysis is conducted on several classical approaches.
Definition 3.1 ().
@@ 3.1. Models of Quantum Complexity Growth
The change of states of a closed quantum mechanical system over time is caused by a specific class of transformations, mathematically known as unitary transformations. Denote the quantum complexity of a unitary transformation or quantum state as the size of the shortest quantum computation that runs the unitary or prepares the state.
One might find it reasonable to expect that the complexity of a quantum state governed by a chaotic many-body Hamiltonian grows linearly with time s.t. the time is exponential in the system size. However, it is uneasy to rule out a shortcut that improves computation efficiency. So, it is hard to derive lower bounds on quantum complexity for particular unitaries, or states without making additional assumptions.
Brandao et al. (brandao2021models, ) establish a growth of the quantum complexity in the time evolution of several models. They prove that with overwhelming probability, an element sampled from an approximate unitary $k$ design has a strong complexity that scales at least linearly in $k$. Unitary k-designs is a framework that lets one approximate randomization over the unitary group with a finite set of unitaries. Without using less than $k+1$ copies of unitaries from these distributions, the approximation cannot be distinguished from the random distribution.
Brandao et al. (brandao2021models, ) also state that using the known relations between the evolution time and the design order $k$ establishes a lower bound on the growth of quantum complexity. For random quantum circuits in particular, the authors make substantial progress on conjectures by Brown and Susskind (PhysRevD.97.086015, ). In addition, by using an established linear relation between the circuit size and design order, they prove a linear growth of quantum complexity.
@@ 3.2. Quantum Randomness and Classical Complexity
Arora et al. (arora2023quantum, ) suggested a comprehensive characterization of the computational power of shallow quantum circuits. Such circuits are combined with classical computation. In particular, for classes of search problems, the authors proved that relative to a random oracle, several properties hold.
First, the authors proved that the conjecture of Jozsa (jozsa2006introduction, ) has to be refused. Becuase $\text{BPP}^{\text{QNC}^{\text{BPP}}}\neq\text{BQP}$. Consequently, an instantiatable separation between the classes was obtained in (arora2023quantum, ). This was done by replacing the oracle with a cryptographic hash function. The latter also implied a resolution to one of Aaronson’s ten semi-grand challenges (aaronson2005ten, ) in quantum computing.
Second, the authors showed that $\text{BPP}^{\text{QNC}}\not\subseteq\text{QNC}^{\text{BPP}}$. Also, they proved $\text{QNC}^{\text{BPP}}\not\subseteq\text{BPP}^{\text{QNC}}$ holds. So, one may observe that an interaction between the classical and the shallow quantum computation exists. In particular, for some problems, performing adaptive measurements in a single shallow quantum circuit sounds more favorable than performing polynomially many shallow quantum circuits, having no adaptive measurements.
Third, the authors stated that there is a 2-message proof of quantum depth protocol. One may instantiate the proof of quantum depth using the one for quantumness construction discussed in the work of Yamakawa and Zhandry (yamakawa2022verifiable, ). One may observe that the above protocol lets a classical verifier certify that a prover has to perform a computation of some minimum quantum depth.
Clifford and Clifford (clifford2018classical, ) studied the classical complexity of the exact Boson Sampling problem where the objective is to produce provably correct random samples from a particular quantum mechanical distribution. Aaronson and Arkhipov proposed the computational framework in STOC ’11 in 2011 as an attainable demonstration of ‘quantum supremacy’, that is a practical quantum computing experiment able to produce output at a speed beyond the reach of classical (that is non-quantum) computer hardware. Aaronson and Arkhipov demonstrated that exact Boson Sampling is not efficiently solvable by a classical computer unless $P^{\#P}=BPP^{NP}$ and the polynomial hierarchy collapses to the third level (aaronson2010computational, ).
Clifford and Clifford give an algorithm running in $\mathcal{O}(n2^{n}+poly(m,n))$ time and $\mathcal{O}(m)$ space. The algorithm is simple to implement and has low constant factor overheads which able it to solve the exact Boson Sampling problem for system sizes far beyond current photonic quantum computing experimentation (clifford2018classical, ).
@@ 3.3. Approximation for Quantum Problems
Gharibian and Kempe (gharibian2012hardness, ) defined a quantum generalization of the polynomial hierarchy and initiated its study. They showed that there are complete problems for the second level of this quantum hierarchy. Further, these problems are hard to approximate. The authors also obtained the hardness of approximation for the class QMA using these techniques. Inspired by the classical results of Umans on the hardness of approximation for the second level of the classical polynomial hierarchy, the authors’ approach is based on the use of dispersers. They proved the hardness of approximation of the problems, including a quantum version of the Succinct Set Cover problem. Further, for a variant of the local Hamiltonian problem with hybrid classical-quantum ground states (gharibian2012hardness, ).
A crucial element in computation is by including quantumness or quantum correlations. These are lasting in entanglements yet often separable states. Consequently, quantifying the quantumness of a state in a quantum system indeed brings useful insights. Identifying the nearest classical-classical state helps in understanding the degree of quantum correlation (lu2023approximation, ). Nevertheless, the exact quantification of quantumness is an NP-hard problem. So, Lu et al. (lu2023approximation, ) considered approximating the exact solution. They took the Frobenius norm, constructed an objective function, and offered a gradient-driven descent flow on Stiefel manifolds. This could determine the quantity. The authors proved that the objective value decreases along the flow. They also provided supporting experimental results. Further, the ability to decompose quantum states into tensor products of certain structures is guaranteed by their method. The method also maintains primitive quantum assumptions. Moreover, the provided numerical results verify that their method is practice-friendly as well.
@@
Let $y$ be a real random variable. Consider $n$ independent samples $y_{1},y_{2},\cdots,y_{n}$. One might estimate the mean of $y$ i.e. $\mu=E[y]$ from the samples. A straightforward approach is to output the mean of the samples: $\hat{\mu}=(y_{1}+\cdots+y_{n})/n$. This is an unbiased estimator with standard deviation $\sigma/\sqrt{n}$ s.t. $\sigma=\textbf{stddev}[y]=E[(y-\mu)^{2}]$. One may observe that Chebyshev’s inequality implies:
@@
Cornelissen et al. (cornelissen2022near, ) proposed a near-optimal quantum algorithm for estimating the mean of a vector-valued random variable with finite mean and covariance, in Euclidean norm. The authors considered the binary oracle model. This generalizes in the classical sample complexity. It also is a setting used frequently in previous works, e.g. (abrams1999fast, ; brassard2011optimal, ; brassard2002quantum, ; hamoudi:LIPIcs.ESA.2021.50, ; HEINRICH20021, ; montanaro2015quantum, ). Cornelissen et al. extended the theory of multivariate sub-Gaussian estimators (lugosi2019mean, ) to the quantum setting. The authors pointed out that under the quantum settings, one may not prove that there is a univariate estimator that can be utilized to become a multivariate estimator, with at most a logarithmic overhead in the dimension. If the sample complexity is considered smaller than the dimension, there exists a quantum advantage for the mean estimation (heinrich2004power, ). Denote $d$ as the dimensionality factor. The lower bound proved in (heinrich2004power, ) implies that a $\log(d)$ overhead for the quantum multivariate mean estimation problem is possible. The crucial point shown in (cornelissen2022near, ) is that apart from the above settings, there is a quantum estimator, outperforming all classical ones. In particular, the quantum estimator reduces the error by a factor close to the square root of the ratio between the dimensionality factor, and the complexity of the sample.
Cornelissen and Hamoudi (cornelissen2023sublinear, ) introduced a quantum algorithm, estimating Gibbs partition functions in sublinear time w.r.t. the logarithm of the size of the state space. The authors obtained a speed-up over the work of Štefankovič et al. (vstefankovivc2009adaptive, ). They utilized the properties of quantum Markov chains to conserve the quadratic speed-up in precision and spectral gap, obtained in previous works. The authors offered polynomial improvements to compute the partition function of the Ising model, counting the number of k-colorings, matching, or independent sets of a graph, and estimating the volume of a convex body. The approach in (cornelissen2023sublinear, ) included developing variants of the quantum phase and amplitude estimation algorithms. To reduce the variance quadratically faster compared to the classical empirical mean, the authors extended the latter subroutines into a nearly unbiased quantum mean estimator.
@@
The Tail Assignment problem is the task of assigning individual aircraft to a given set of flights, minimizing the overall cost. Vikstål et al. (vikstaal2020applying, ) simulated the QAOA applied to instances of this problem derived from real-world data of the Tail Assignment problem. These instances were reduced to fit on quantum devices with 8, 15, and 25 qubits. For instance, the reduction process leaves only one viable solution that allows us to map out an Exact Cover problem on Tail Assignment (vikstaal2020applying, ).

### v4

@@ 4. Quantum Algorithms Complexity
Keeping in mind the struggle to connect classical computation and quantum in mind, consider a random oracle. Yamakawa and Zhandry (yamakawa2022verifiable, ) proved unconditionally unless otherwise stated, that relative to the oracle, with probability one, it holds that:
@@
  • The results of Yamakawa and Zhandry do not contradict the conjecture of Aaronson and Ambanis (aaronson2014need, ). As a conjecture, they mentioned that considering the Aaronson-Ambanis conjecture is true, one may conclude that there is publicly verifiable certifiable randomness with minimal rounds of interaction.
@@
Away from the BQP, yet another property of quantum computation was exploited in the literature. It is known that the structure of optimally entangled states is constrained by quantum theory’s constant character (alber1999entanglement, ). Abrams et al. (abrams1998nonlinear, ) illustrated that one may exploit nonlinear time evolution to allow a quantum computer to solve NP-complete and #P problems in polynomial time. The authors showed how to accomplish this exponential speed-up using the Weinberg model of nonlinear quantum mechanics. All known experiments have confirmed the linearity of quantum mechanics to a high degree of accuracy. So, the result in (abrams1998nonlinear, ) may be seen as evidence.
@@
Interestingly, an entirely different approach is somewhat helping the quantum side with the classical. Wang (wang2022classically, ) leveraged existing classical techniques to improve quantum optimization. In particular, he first pre-processed the input with a classical algorithm, obtaining an approximate solution. Second, a quantum circuit to search the neighbourhood of the solution of the classical algorithm, aiming for higher-quality solutions. They introduced the Classically-Boosted Quantum Optimization Algorithm (CBQOA). One may solve a wide range of combinatorial optimization problems with CBQOA. E.g. all unconstrained problems, constrained problems such as Max Bisection, Maximum Independent Set, Minimum Vertex Cover, Portfolio Optimization, Traveling Salesperson, etc. A crucial component of CBQOA is a practice-friendly continuous-time quantum walk (CTQW) on a properly constructed graph, connecting the feasible solutions. CBQOA utilizes CTQW, the output of a classical procedure, constructing a superposition of feasible solutions. The algorithm solves constrained problems without applying changes to their cost functions. It then narrows down evaluating the quantum state to the feasible subspace. The algorithm also relies not on indexing the feasible solutions. Wang illustrates the applications of CBQOA to Max 3SAT, and Max Bisection as well. The author further provides empirical results that CBQOA outperforms previous approaches.
@@
Apart from the above approach and turning to the prover-verifier model, consider the Quantum Merlin Arthur (QMA). That represents the set of languages for which, when a string is in the language, there is a polynomial-size quantum proof that convinces a polynomial time quantum verifier of this fact with high probability. In addition, the verifier with high probability rejects every linear size quantum state where no string is included in the language.
Some methods suggest using the class $\text{QMA}^{+}(2)$, that is, the unentangled quantum proofs with non-negative amplitudes. Jeronimo and Wu (10.1145/3564246.3585248, ) showed that under such settings, one may design proof verification protocols for (increasingly) hard problems. This is done via logarithmic size quantum proofs and also by having a constant probability gap that allows distinguishing yes from no instances. Specifically, the authors designed global protocols for small set expansion, unique games, and PCP verification. Consequently, Jeronimo and Wu had the result of $\text{NP}\subseteq\text{QMA}^{+}_{\log}(2)$ with a constant gap. Taking advantage of the mentioned constant gap, the authors could scale up the aforementioned result to $\text{QMA}^{+}(2)$. So, they were able to obtain $\text{QMA}^{+}(2)=\text{NEXP}$ by setting up stronger in-detail properties of the PCP for NEXP.
Gottesman and Irani (gottesman2009quantum, ) have studied the complexity of problems that involve satisfying constraints remaining the same under translations in at least one spatial direction. Given an input $N$ in binary. The authors proved the hardness of a classical tiling problem on an $N\crossproduct N$ 2-dimensional grid; And a quantum problem involving finding the ground state energy of a 1-dimensional quantum system of $N$ particles. The classical problem is then proved to be NEXP-complete and the quantum problem is ${\textit{QMA}}_{\mathit{E}XP}$ complete. Thus, an algorithm for these problems that runs in time polynomial in $N$ would imply that $\mathit{E}XP=$ NEXP or BQEXP$=\textit{QMA}_{\mathit{E}XP}$, respectively. Tiling is known to be NEXP-complete. Nevertheless, all previous reductions require that $(1)$ the set of tiles and their constraints $(2)$ the varying boundary conditions, be given as part of the input. Gottesman and Irani considered the problem instance encoded solely in the size of the system.
@@
There are even more interesting results about the QMA complexity class. One may generalize the satisfiability problem in classical computation to a constraint satisfaction problem in the quantum model, called Quantum Satisfiability. Consider a quantum k-SAT problem. Each constraint is then determined by a k-local projector. Also, a constraint is satisfied by all states in its null space. A previous work by Bravyi (bravyi2011efficient, ) showed that one may solve quantum 2-SAT on a classical computer. The author further proved that quantum k-SAT with $k\geq 4$ is QMA-complete. Previously, Quantum 3-SAT was recognized as being in QMA (bravyi2011efficient, ). Nonetheless, its computational hardness was unknown. Gosset and Nagaj (gosset2016quantum, ) then proved that quantum 3-SAT is QMA-hard. So, this problem is complete for this complexity class.
Another direction in the literature is exploiting the properties of quantum entanglement to help improve the known classical complexity of certain problems. More specifically, consider the classical matching theory, that is, a mathematical method for finding pairs between elements from two sets. One may optimize it according to certain criteria. Classical matching theory can be defined as matrices that do not contain negative entries. In quantum theory, the positive operator concept is a generalization of matrices that have nonnegative entries. In terms of complete positive operators, or equivalently with bipartite density matrices, Gurvits (GURVITS2004448, ) reformulate the Edmonds problem. Namely, the central problem in the field of algebraic complexity, which J. Edmonds presented in 1967. This problem aims to decide if the span of a given tuple of square matrices contains a non-singular matrix. In other words, it investigates whether a given linear subspace of a matrix has a full rank. One of the most important cases when Edmonds’ problem can be solved in polynomial deterministic time, i.e., an intersection of two geometric matroids, corresponds to unentangled (a.k.a. separable) bipartite density matrices.
Denote $M(N)$ as the linear space of complex $N\crossproduct N$ matrices. To solve Edmonds’ problem, Gurvits introduced a general class of linear subspaces $M(N)$. In this class, there is a polynomial deterministic time algorithm to solve Edmonds’ problem. Finally, Gurvits proved that the weak membership problem for the convex set of separable normalized bipartite density matrices is NP-hard.
@@
Khadiev and Safina (khadiev2023quantum, ) suggested a quantum algorithm for the dynamic programming approach for problems on directed acyclic graphs (DAGs) like $G(V,E)$. Denote $|V|=n$ as the number of vertices of $G(V,E)$. Further, let $\hat{n}$ be the number of vertices with at least one outgoing edge, and $|E|=m$ be the number of edges. Their algorithm’s running time is proportional to $\mathcal{O}(\sqrt{\hat{n}m}\log\hat{n})$. Moreover, executing the best-known deterministic algorithm up to the date of the publication of (khadiev2023quantum, ) takes $\mathcal{O}(n+m)$. The authors proved that their algorithm may solve problems that involve functions OR, AND, NAND, MAX, and MIN. For example, a Boolean formula represented by Zhegalkin polynomial (zhegalkin1927technique, ; gindikin1985algebraic, ; yablonsky1989introduction, ). Also, a Boolean circuit with shared input and non-constant depth evaluation. In addition, the single source longest paths search for weighted DAGs problem and the diameter search problem for unweighted DAGs.
@@
Moving on from the above directions, consider the Quantum Annealing algorithms, introduced as a solution to combinatorial optimization problems. They utilise quantum fluctuations to escape local minima in energy landscapes to solve complex NP-hard problems. Cruz-Santos et al. (cruz2019qubo, ) aimed to apply quantum annealing to the theory of cuts, which is a crucial field in theoretical computer science. The authors developed a method to convert the Minimum Multicut Problem into the Quadratic Unconstrained Binary Optimization (QUBO) representation and have addressed the technical challenges that arise when submitting a problem to the quantum annealer processor. Two constructions of the quadratic unconstrained binary optimization functions for the Minimum Multicut Problem have been provided and compared the tradeoffs between the two mappings. Also, numerical scaling analysis is conducted on several classical approaches.
Definition 4.1.
@@ 4.1. Models of Quantum Complexity Growth
The change of states of a closed quantum mechanical system over time is caused by a specific class of transformations, mathematically known as unitary transformations. Denote the quantum complexity of a unitary transformation or quantum state as the size of the shortest quantum computation that runs the unitary or prepares the state. One might find it reasonable to expect that the complexity of a quantum state governed by a chaotic many-body Hamiltonian grows linearly with time s.t. the time is exponential in the system size. However, it is uneasy to rule out a shortcut that improves computation efficiency. So, it is hard to derive lower bounds on quantum complexity for particular unitaries, or states without making additional assumptions.
Brandao et al. (brandao2021models, ) establish a growth of the quantum complexity in the time evolution of several models. They prove that with overwhelming probability, an element sampled from an approximate unitary $k$ design has a strong complexity that scales at least linearly in $k$. Unitary k-designs is a framework that lets one approximate randomization over the unitary group with a finite set of unitaries. Without using less than $k+1$ copies of unitaries from these distributions, the approximation cannot be distinguished from the random distribution. They also state that using the known relations between the evolution time and the design order $k$ establishes a lower bound on the growth of quantum complexity. For random quantum circuits in particular, the authors make substantial progress on conjectures by Brown and Susskind (PhysRevD.97.086015, ). In addition, by using an established linear relation between the circuit size and design order, they prove a linear growth of quantum complexity.
@@ 4.2. Quantum Randomness and Classical Complexity
Arora et al. (arora2023quantum, ) suggested a comprehensive characterization of the computational power of shallow quantum circuits. Such circuits are combined with classical computation. In particular, for classes of search problems, the authors proved that relative to a random oracle, several properties hold. First, they proved that the conjecture of Jozsa (jozsa2006introduction, ) has to be refused. Becuase $\text{BPP}^{\text{QNC}^{\text{BPP}}}\neq\text{BQP}$. Consequently, an instantiatable separation between the classes was obtained in (arora2023quantum, ). This was done by replacing the oracle with a cryptographic hash function. The latter also implied a resolution to one of Aaronson’s ten semi-grand challenges (aaronson2005ten, ) in quantum computing. Second, they showed that $\text{BPP}^{\text{QNC}}\not\subseteq\text{QNC}^{\text{BPP}}$. Also, they proved $\text{QNC}^{\text{BPP}}\not\subseteq\text{BPP}^{\text{QNC}}$ holds. So, one may observe that an interaction between the classical and the shallow quantum computation exists. In particular, for some problems, performing adaptive measurements in a single shallow quantum circuit sounds more favourable than performing polynomially many shallow quantum circuits, having no adaptive measurements. Third, they stated that there is a 2-message proof of quantum depth protocol. One may instantiate the proof of quantum depth using the one for quantumness construction previously discussed in the work of Yamakawa and Zhandry (yamakawa2022verifiable, ). One may observe that the above protocol lets a classical verifier certify that a prover has to perform a computation of some minimum quantum depth.
Clifford and Clifford (clifford2018classical, ) studied the classical complexity of the exact Boson Sampling problem where the objective is to produce provably correct random samples from a particular quantum mechanical distribution. Aaronson and Arkhipov proposed the computational framework in STOC ’11 in 2011 as an attainable demonstration of ‘quantum supremacy’, that is a practical quantum computing experiment able to produce output at a speed beyond the reach of classical (that is non-quantum) computer hardware. Aaronson and Arkhipov demonstrated that exact Boson Sampling is not efficiently solvable by a classical computer unless $P^{\#P}=BPP^{NP}$ and the polynomial hierarchy collapses to the third level (aaronson2010computational, ). Clifford and Clifford give an algorithm running in $\mathcal{O}(n2^{n}+poly(m,n))$ time and $\mathcal{O}(m)$ space. The algorithm is simple to implement and has low constant factor overheads which able it to solve the exact Boson Sampling problem for system sizes far beyond current photonic quantum computing experimentation (clifford2018classical, ).
@@ 4.3. Approximation for Quantum Problems
Gharibian and Kempe (gharibian2012hardness, ) defined a quantum generalization of the polynomial hierarchy and initiated its study. They showed that there are complete problems for the second level of this quantum hierarchy. Further, these problems are hard to approximate. They further obtained the hardness of approximation for the class QMA using these techniques. Inspired by the classical results of Umans on the hardness of approximation for the second level of the classical polynomial hierarchy, the authors’ approach is based on the use of dispersers. They proved the hardness of the approximation of the problems, including a quantum version of the Succinct Set Cover problem. Further, for a variant of the local Hamiltonian problem with hybrid classical-quantum ground states (gharibian2012hardness, ).
A crucial element in computation is by including quantumness or quantum correlations. These are lasting in entanglements yet often separable states. Consequently, quantifying the quantumness of a state in a quantum system indeed brings useful insights. Identifying the nearest classical-classical state helps in understanding the degree of quantum correlation (lu2023approximation, ). Nevertheless, the exact quantification of quantumness is an NP-hard problem.
Lu et al. (lu2023approximation, ) considered approximating the exact solution. They took the Frobenius norm, constructed an objective function, and offered a gradient-driven descent flow on Stiefel manifolds. This could determine the quantity. The authors proved that the objective value decreases along the flow. They also provided supporting experimental results. Further, the ability to decompose quantum states into tensor products of certain structures is guaranteed by their method. The method also maintains primitive quantum assumptions. Moreover, the provided numerical results verify that their method is practice-friendly as well.
@@
Moving on from QMC, let $y$ be a real random variable. Consider $n$ independent samples $y_{1},y_{2},\cdots,y_{n}$. One might estimate the mean of $y$ i.e. $\mu=E[y]$ from the samples. A straightforward approach is to output the mean of the samples: $\hat{\mu}=(y_{1}+\cdots+y_{n})/n$. This is an unbiased estimator with standard deviation $\sigma/\sqrt{n}$ s.t. $\sigma=\textbf{stddev}[y]=E[(y-\mu)^{2}]$. One may observe that Chebyshev’s inequality implies:
@@
Another interesting result is obtained by Cornelissen et al. (cornelissen2022near, ). They proposed a near-optimal quantum algorithm for estimating the mean of a vector-valued random variable with finite mean and covariance, in Euclidean norm. The authors considered the binary oracle model. This generalizes in the classical sample complexity. It also is a setting used frequently in previous works, e.g. (abrams1999fast, ; brassard2011optimal, ; brassard2002quantum, ; hamoudi:LIPIcs.ESA.2021.50, ; HEINRICH20021, ; montanaro2015quantum, ). Cornelissen et al. extended the theory of multivariate sub-Gaussian estimators (lugosi2019mean, ) to the quantum setting. The authors pointed out that under the quantum settings, one may not prove that there is a univariate estimator that can be utilized to become a multivariate estimator, with at most a logarithmic overhead in the dimension. If the sample complexity is considered smaller than the dimension, there exists a quantum advantage for the mean estimation (heinrich2004power, ). Denote $d$ as the dimensionality factor. The lower bound proved in (heinrich2004power, ) implies that a $\log(d)$ overhead for the quantum multivariate mean estimation problem is possible. The crucial point shown in (cornelissen2022near, ) is that apart from the above settings, there is a quantum estimator, outperforming all classical ones. In particular, the quantum estimator reduces the error by a factor close to the square root of the ratio between the dimensionality factor, and the complexity of the sample.
Cornelissen and Hamoudi (cornelissen2023sublinear, ) introduced a quantum algorithm, estimating Gibbs partition functions in sublinear time w.r.t. the logarithm of the size of the state space. The authors obtained a speed-up over the work of Štefankovič et al. (vstefankovivc2009adaptive, ). They utilized the properties of quantum Markov chains to conserve the quadratic speed-up in precision and spectral gap, obtained in previous works. The authors offered polynomial improvements to compute the partition function of the Ising model, counting the number of k-colourings, matching, or independent sets of a graph, and estimating the volume of a convex body. The approach in (cornelissen2023sublinear, ) included developing variants of the quantum phase and amplitude estimation algorithms. To reduce the variance quadratically faster compared to the classical empirical mean, the authors extended the latter subroutines into a nearly unbiased quantum mean estimator.
Definition 4.2.
The Tail Assignment problem is the task of assigning individual aircraft to a given set of flights, minimizing the overall cost.
Definition 4.3.
@@
Vikstål et al. (vikstaal2020applying, ) simulated the QAOA applied to instances of this problem derived from real-world data of the Tail Assignment problem. These instances were reduced to fit on quantum devices with 8, 15, and 25 qubits. For instance, the reduction process leaves only one viable solution that allows us to map out an Exact Cover problem on Tail Assignment (vikstaal2020applying, ).